OpenStudy (kasper-kay):

A line passes through the point (–2, 4), and its y-intercept is (0, –6). What is the equation of the line that is perpendicular to the first line and passes through the point (5, –4)?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

first find the slope do you know how to do that?

OpenStudy (kasper-kay):

y- y1 = m (x-x1)

OpenStudy (misty1212):

no dear that is the point - slope formula you need the slope first, then you can use that one

OpenStudy (kasper-kay):

then i dont know how to find the slope

OpenStudy (misty1212):

\[\frac{y_2-y_1}{x_2-x_1}\] with \[x_1=-2,y_1=4,x_2=0,y_2=-6\]

OpenStudy (kasper-kay):

so the slope is -5/1

OpenStudy (misty1212):

or you can reason as follows from -2 to 0 is right 2 units from 4 to -6 is down ten right 2, down 10 slope is \(-\frac{10}{5}=-5\)

OpenStudy (misty1212):

now you need the slope of the perpendicular line that will be the "negative reciprocal" of \(-5\) which is \(\frac{1}{5}\)

OpenStudy (misty1212):

NOW you can use the point slope formula with the slope \(m=\frac{1}{5}\) and the point \((5,-4)\)

OpenStudy (kasper-kay):

so y- -4 = 1/5 (x - 5)?

OpenStudy (kasper-kay):

So the anser would be y = 1/5x - 5